Jenny Is Older Than Paul The Difference Of Their Ages Is Eight And The Sum Of Their age Is Sixty Find The Age Of Each
Answer by Maths68(1474)

Let
Present age of Jenny = j
Present age of Paul = p
Jenny Is Older Than Paul The Difference Of Their Ages Is Eight
j-p=8..........(1)
The Sum Of Their age Is Sixty
j+p=60.........(2)
Add (1) and (2)
j-p=8..........(1)
j+p=60.........(2)
--------
2j=68
2j/2=68/2
j=34
Put the value of j in (2)
j+p=60.........(2)
34+p=60
p=60-34
p=26

Present age of Jenny = j=34 years old
Present age of Paul = p = 26 years old